OpenStudy (anonymous):

True, False or Open 9p+8 = 10p + 7 helpp?

OpenStudy (anonymous):

p= 1

OpenStudy (anonymous):

true only when p=1

OpenStudy (anonymous):

9p+8 = 10p + 7 First lets solve. Subtract 9p from both sides -9p -9p --------------- 8 = 1p + 7 subtract 7 from both sides -7 -7 ------------- 1 = 1p Divide /1 /1 ----------- 1 = p True :)

OpenStudy (phi):

True, False or Open ? It is not always True, and it is not always False. It depends on what "p" is. What do you think?